OpenStudy (anonymous):

A square piece of paper is folded in half the perimeter Of the resulting rectangle is 24 cm find the area of the original square paper

OpenStudy (anonymous):

Let the side of square be x When the paper is folded the resultant rectangle will have length same as side of the square and the width 1/2 of the square side Perimeter of rectangle = x+x+x/2+x/2 = 3x 3x=24 x=8 The area of originalaper is 8*8 = 64 square cm